OpenStudy (anonymous):

explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences.

OpenStudy (anonymous):

(2x+3)(x+5)

hero (hero):

\[ac = (\frac{b}{2}+d)(\frac{b}{2}-d)\]

hero (hero):

solve for d

OpenStudy (anonymous):

(2x+3)(x+5)

hero (hero):

\[30 = (\frac{13}{2}+d)(\frac{13}{2}-d)\]

hero (hero):

\[30 = \frac{169}{4}-d^2\]

hero (hero):

\[d^2 = \frac{169}{4} - 30\]

hero (hero):

\[d^2 = \frac{169}{4}-\frac{120}{4}\]

hero (hero):

\[d = \sqrt{\frac{49}{4}}\]

hero (hero):

\[d = \frac{7}{2}\]

hero (hero):

\[(\frac{13}{2} + \frac{7}{2})(\frac{13}{2}-\frac{7}{2})\]

hero (hero):

\[(10)(3)\]

hero (hero):

Those are your two middle numbers

OpenStudy (anonymous):

thanks 4 being so detailed

hero (hero):

So now, write it like this: \[2x^2+10x + 3x + 15\]

hero (hero):

Then factor out what's common to the first two terms and what's common to the last two terms

hero (hero):

\[2x(x+5)+3(x+5) = (x+5)(2x+3)\]

hero (hero):

You should be able to make plenty sentences out of that

hero (hero):

Make sure you start off saying to set \[a = (\frac{b}{2}+d)\] and \[b = (\frac{b}{2}-d)\]

hero (hero):

Your teacher will probably be intrigued by all this.